# Completed Feature Disentanglement Learning for Multimodal MRIs Analysis

### v1

@@ Abstract
Figure 1: Illustrations of incompleted feature representation for existing feature disentanglement methods in the three-modal case. Sub-figure (a) presents a concept map, (b) shows a case in the Meningiomas dataset. In sub-figure (b), the tumor and edema areas are enclosed within the red and the green boxes respectively, and the yellow box indicates high cell density in tumor area. Notably, the location information of tumor and edema is shared across all three MRI series. Additionally, the tumor characteristic information is shared between the T1C and FLAIR-C series, the edema characteristics are shared between the FLAIR-C and ADC series, while the information regarding the cell density in tumor area is shared between the T1C and ADC series.
@@ 1 Introduction
Multi-modality data contains multiple aspects of information about an object, and different modalities can provide complementary information. Numerous previous studies have demonstrated the remarkable success of multimodal learning (MML) [1] for medical image analysis. However, inappropriate processing of multimodal information can significantly impact the efficiency of MML. According to [2], the key to successful MML lies in achieving a higher quality of feature representation. Many previous works [3, 4, 5, 6, 7, 8, 9, 10, 11] have focused on enhancing the learning performance, which can be classified into three categories. Two categories focus on extracting the shared information between multiple modalities [4, 3] or specific information of each modality [5, 6]. These methods cannot fully extract multimodal information, as they only focus on one type of feature, leading to information loss [11, 10]. The third category focuses on feature disentanglement (FD), which simultaneously decouples modality-shared features as well as modality-specific features [7, 9, 8, 10], leading to sound results. These FD-based methods commonly enforce the following constraints on decoupled features: 1) modality-shared features should have higher similarity; 2) modality-specific features should have greater dissimilarity; and 3) modality-shared features and each modality-specific features should be as distinct as possible.
@@
We revisit the relationship between the representation spaces of multimodal data. As illustrated in the concept map of the three-modal case in Fig. 1, intuitively, we can consider that there exists modality-shared information (yellow area) as well as modality-specific information (green area). Upon further consideration and exploration, we discover that there is shared information present between subsets of modalities (blue area). However, existing FD methods could potentially ignore such information due to the above three constraints, which only account for modality-shared information and modality-specific information. On the other hand, our preliminary experiments reveals that the lost information is crucial for accurate prediction (see the first and fourth rows of Table 4). In Fig. 1 (b), we take meningiomas grading and invasion prediction data as an example for illustration. We can see that different MRI series provide specific information based on their distinct signal characteristics. The tumor area is highlighted in both T1C and FLAIR-C, indicating that these two modalities share information about the tumor. The edema area is highlighted in both FLAIR-C and ADC, indicating that both MRI series offer information about the edema characteristics. Additionally, T1C and ADC share information about the cell density in tumor area. In fact, such shared information among pair-wise modalities is found to be relevant to the prediction of meningiomas grade and invasion in clinical research [12, 13, 14].
Furthermore, existing FD-based approaches also struggle with determining the optimal fusion methods for decoupled features. Current fusion studies mainly focus on uncertainty-based fusion methods [15, 16] and attention-based methods [17, 18]. Uncertainty-based methods may introduce more uncertainty when dealing with difficult-to-predict classes, potentially resulting in incorrect predictions. Attention-based methods perform feature fusion implicitly and cannot explicitly reveal the relationship between multiple features.
@@
To tackle the above issues, we propose a completed feature disentanglement multimodal learning (CFDL) approach for multimodal MRIs analysis. First, we present a novel completed feature disentanglement (CFD) strategy to address the information loss in previous FD-based methods. In addition to decoupling modality-shared features among all modalities and modality-specific features, we further decouple features shared between subsets of modalities, referred to as modality-partial-shared features. The modality partial-shared features are also expected to have higher similarities, while to be dissimilar from the other two kinds of features. To improve the interpretability of feature fusion, we propose a new dynamic mixture-of-experts fusion (DMF) module, which can explictly capture local-global interrelationships between the decoupled features for better fusion. In summary, our main contributions are summarized as follows:
  • We find that existing FD-based MML methods suffer from the information loss due to incomplete feature decoupling and the lack of interpretability for the decoupled features fusion.
  • We propose a novel CFD strategy to further decouple modality-partial-shared features to ensure completed feature decoupling from multimodal data. We demonstrate that these features play a crucial role in prediction.
  • We introduce a new DMF module which can dynamically integrate the decoupled features while explicitly revealing the relationships between these features. This module enhances the interpretability of the fusion process.
  • Experiments on three multimodal MRI datasets demonstrate the effectiveness and superiority of our framework compared to the state-of-the-art methods.
@@ 2.2 Multimodal Feature Fusion
In recent years, there has been an increasing focus on exploring the correlations among multiple modalities to obtain effective features. Some methods achieve multimodal fusion by assigning weights or probabilities to each modality. Han et al. [15, 16] explored the uncertainties of different modalities to obtain reliable multimodal fusion information. Choi et al. [20] proposed EmbraceNet which performed multimodal representation fusion based on a probabilistic approach. However, these methods could be subject to the effect of difficult-to-predict classes. Zhou et al. [21] introduced a canonical correlation analysis (CCA)-based method named ADCCA to exploit the correlation between multiple modalities and integrate the complementary information from these modalities.
@@ 3.1 Overview
The proposed CFDL framework, as shown in Fig. 2, consists of three parts: a) feature extraction from multimodal MRIs, b) completed feature disentanglement for feature decoupling, and c) dynamic MoE fusion for dynamically integrating the decoupled features. The framework employs the same type of backbone to extract latent features $x_{j}$ from each modality. To capture a comprehensive representation of the multimodal data, the latent features are decoupled into modality-shared features, modality-specific features, and modality-partial-shared features using the proposed CFD strategy. The decoupled features are then integrated using the DMF module. In the DMF module, each decoupled feature is paired with a specific expert, and a gating network named LinG_GN generates weights for multiple experts. The fused features are obtained by aggregating the weighted features from the experts. In the following, we take three-modal condition as examples to illustrate the proposed CFD strategy and DMF module.
@@ 3.2 Completed Feature Disentanglement Strategy
Inspired by previous FD methods [7, 8, 9], we first decouple the extracted latent features into modality-shared features and modality-specific features. We employ a shared-encoder $\bm{E_{sh}}$ to decouple modality-shared features and three private-encoders $\bm{E_{sp}}^{j}$ to decouple modality-specific feature for each modality. Three modality-shared features can be formulated as follows:
$F_{j}=\bm{E_{sh}}(x_{j}),$ (1)
@@
We further consider modality-partial-shared features between pair-wise modalities. As a result, three groups of modality-partial-shared features are decoupled, with each group consisting of two features. The two features in the same group ($G_{jk}^{j}$, $G_{jk}^{k}$) are decoupled with the same partial-shared-encoder named $\bm{E_{ps}}^{jk}$, i.e.,
$G_{jk}^{j}=\bm{E_{ps}}^{jk}(x_{j}),\quad G_{jk}^{k}=\bm{E_{ps}}^{jk}(x_{k}).$ (4)
@@
To ensure the effective decoupling of modality-shared features and modality-partial-shared features, we employ the mean squared error (MSE) loss as a constraint. The MSE loss measures the discrepancy between two features, and we aim to increase similarity between two features by minimizing this loss. The losses for modality-shared features $\mathcal{L}_{sh}$ and modality-partial-shared features $\mathcal{L}_{ps}$ are expressed as:
$\mathcal{L}_{sh}=\sum_{j=1}\sum_{k=j+1}MSE(F_{j},F_{k}),$ (6)
$\mathcal{L}_{ps}=\sum_{j=1}\sum_{k=j+1}MSE(G_{jk}^{j},G_{jk}^{k}).$ (7)
@@ 3.3 Dynamic MoE Fusion Module
To ensure dynamic fusion of the final decoupled features, we introduce the DMF module based on the MoE architecture, which is shown in Fig. 2 (c). In DMF module, each final decoupled feature $\mathcal{S}_{j}\in\mathbb{R}^{dim}$ is associated with a specific expert $\bm{Ex}_{j}$, which is implemented as a fully-connected layer. We introduce LinG_GN to dynamically generate weights for these experts, taking into account the relationships of these final decoupled features. The LinG_GN operates with two inputs to capture a comprehensive understanding of these features. Firstly, we concatenate all the final decoupled features as one input for the LinG_GN. To integrate this concatenated feature, we map the concatenated feature into a unified representation utilizing a fully-connected layer $FC$. This unified representation captures the collective information from all the final decoupled features and is treated as the global feature $g\in\mathbb{R}^{dim}$,
@@
where $\times$ represents matrix multiplication operation, $\omega\in\mathbb{R}^{7}$. Each element in $\omega$ represents the weight for the corresponding expert network $\bm{Ex}_{j}$. The fused feature $F_{f}$ is obtained by concatenating the linearly weighted experts,
$F_{f}=Cat(\omega_{1}\bm{Ex}_{1}(\mathcal{S}_{1}),...,\omega_{7}\bm{Ex}_{7}(\mathcal{S}_{7})).$ (11)
@@
$\mathcal{L}=\mathcal{L}_{cls}+\alpha(\mathcal{L}_{sh}+\mathcal{L}_{ps})+\beta\mathcal{L}_{diff},$ (13)
@@ 3.4 Network Architecture
We utilize the 3D ResNet18 [28] as the backbone for feature extraction, and the parameters of these backbones are not shared. The dimension of the each extracted latent feature $x_{j}$ is 512. The shared-encoder $\bm{E_{sh}}$, each private-encoders $\bm{E_{sp}}^{j}$, each partial-shared-encoders $\bm{E_{ps}}^{jk}$, $FC$ and each expert $\bm{Ex}_{j}$ are all implemented as one fully-connected layer with $dim$ neurons. The $MLP$ consists of two fully-connected layers with $dim$ neurons and one output layer with $num\_cls$ neurons, where $num\_cls$ represents the number of classes. Each fully-connected layer in $MLP$ is followed by a ReLU layer and a Dropout layer. We empirically set $dim$ as 32.
@@ 4.1.1 Meniscal Tear Prediction
For the prediction of meniscal tear, we employ the MRNet dataset [29], which is a publicly available knee multi-MRI dataset [^1] More information about the MRNet dataset can be available in https://stanfordmlgroup.github.io/competitions/mrnet/.. There are 1130 cases in training set and 120 cases in validation set. The training set contains 397 meniscal tear cases and 733 contrast cases, and the validation set contains 52 meniscal tear cases and 68 contrast cases. Each case includes three MRIs: sagittal plane T2-weighted series (T2-sagittal), coronal plane T1-weighted series (T1-coronal) and axial plane PD-weighted series (PD-axial). We resize the MRIs to 24*128*128 as the inputs of the framework.
@@ 4.1.2 Meningiomas Grading Prediction
We collect Meningiomas Grading Prediction Dataset, referred to as MEN, from the Brain Medical Center of Tianjin University, Tianjin Huanhu Hospital  [^2] The Ethical Committee of Tianjin Huanhu Hospital approves scientific research using these MRIs and waives the need for informed patient consent ((Jinhuan) Ethical Review No.(2022-046)).. This dataset consists of three grades of meningiomas: Grade 1 (G1), Grade 2 with invasion (G2inv) and Grade 2 without invasion (G2ninv). The total dataset comprises 798 cases, including 650 Grade_1, 62 Grade_2inv and 86 Grade_2ninv cases. Each case includes three brain MRIs: Contrast-Enhanced T1 series (T1C), Contrast-Enhanced T2 FLAIR series (FLAIR-C) and Apparent Diffusion Coefficient series (ADC). Following previous works [30, 31], we request radiologists to crop the regions of interest (ROIs). To maintain the shape of the tumor and edema regions, the ROIs are zero-padded into squares and resized to dimensions of 24*128*128, which serve as the inputs for the framework.
@@ 4.1.3 MGMT Promoter Status Prediction
The MGMT Promoter Status Prediction Dataset, known as BraTS 2021 [32], is a publicly available multimodal brain MRI dataset. It encompasses cases with MGMT methylated (MGMT+) and unmethylated (MGMT-) status. The dataset comprises 580 available cases333 We drop 5 cases from original 585 cases during pre-processing. There are 3 cases with unexpected issues. The other 2 cases cannot be registered by CaPTk. Get more information in https://www.kaggle.com/c/rsna-miccai-brain-tumor-radiogenomic-classification., with each case containing four modalities: T1, post-contrast T1-weighted (T1Gd), T2-weighted (T2), and T2 Fluid Attenuated Inversion Recovery (T2-FLAIR). Specifically, there are 275 MGMT- cases and 305 MGMT+ cases. Pre-processing for each modality involves image registration and skull-stripping using the Cancer Imaging Phenomics Toolkit (CaPTk) [33]. We crop the ROIs using masks generated by the pretrained Swin UNETR [34]. Finally, we zero-pad the ROIs into squares and resize them to 16*128*128 as the inputs for the proposed framework.
@@ 4.2.1 Training Details
We employ 3-fold cross-validation for private MEN and public BraTS 2021 datasets, and train three times using different seeds with already divided training and validation data for MRNet dataset. During model training, we implement several techniques to prevent overfitting, such as data augmentation, L2 regularization (weight decay) and dropout [35]. Data augmentation techniques include random clip, random crop, gaussian noise and random erasing [36]. The weight decay is set as $41e-$ 4, and the dropout value is set to $0.5$. The network is optimized with the Adam optimizer [37]. We linearly warm up the learning rate from zero to the preset value over $5$ epochs and apply a learning rate decay strategy, reducing the learning rate to $0.8$ after every $5$ epochs. The batch size is set $32$. For MRNet dataset, we initialize the learning rate value as $48e-$ 4, and set the number of epochs to $50$. The balance factors, $\alpha$ and $\beta$, are set to $0.5$ and $0.005$ respectively. For MEN dataset, the learning rate is specified as $45e-$ 4, and number of epochs is fixed as $100$. Both $\alpha$ and $\beta$ are set as $1$. For BraTS 2021 dataset, we preset the learning rate value as $42e-$ 4 and the number of epochs as $50$. We fix both $\alpha$ and $\beta$ as $0.001$. All experiments are conducted with PyTorch on an NVIDIA RTX 3090 GPU.

### v2

@@ 1 Introduction
Figure 1: A concept map for illustrations of incomplete feature representations in existing FD methods in the three-modal case.
Figure 2: An example from the MEN dataset for illustrations of information loss between pair-wise modalities. Region of interest (ROI), including the tumor, edema and their surrounding area, is marked with an white box and used as the input of the model. Tumor and edema areas are marked with red and green boxes, respectively, and the yellow box highlights regions of high cell density in the tumor. Tumor characteristics are shared between T1C and FLAIR-C, edema characteristics between FLAIR-C and ADC, and cell density information between T1C and ADC.
@@
Multi-modality data contains multiple aspects of information about an object, and different modalities can provide complementary information. Numerous previous studies have demonstrated the remarkable success of multimodal learning (MML) [1] for medical image analysis. However, inappropriate processing of multimodal information can significantly impact the efficiency of MML. According to [2], the key to successful MML lies in achieving a higher quality of feature representation. Many previous works [3, 4, 5, 6, 7, 8, 9, 10, 11] have focused on enhancing the learning performance, which can be classified into three categories. Two categories focus on extracting the shared information between multiple modalities [4, 3] or specific information of each modality [5, 6]. These methods cannot fully extract multimodal information, as they only focus on one type of feature, leading to information loss [11, 10]. The third category focuses on feature disentanglement (FD), which decouples modality-shared features as well as modality-specific features [7, 9, 8, 10], leading to sound results.
@@
We revisit the relationship between the representation spaces of multimodal data. As illustrated in the concept map of the three-modal case in Fig. 1, intuitively, we can consider that there exists modality-shared information (yellow area) as well as modality-specific information (green area). Upon further consideration and exploration, we discover that there is shared information present between subsets of modalities (blue area). However, existing FD methods could potentially ignore such information. On the other hand, our preliminary experiments reveal that the lost information is crucial for accurate prediction (see the first and fourth rows of Table 4). In Fig. 2, we take MEN dataset as an example for illustration. Both T1C and FLAIR-C highlight the tumor area, indicating shared tumor information. Similarly, FLAIR-C and ADC highlight the edema area, showing shared edema information. Additionally, T1C and ADC share information about cell density in the tumor area. In fact, such shared information among pair-wise modalities is found to be relevant to the prediction of meningiomas grade and invasion in clinical research [12, 13, 14].
Furthermore, current multimodal fusion studies mainly focus on uncertainty-based fusion methods [15, 16, 17] and attention-based methods [18, 19]. However, these approaches generally address modality-shared and modality-specific information fusion across multiple modalities, while overlooking the design of fusion mechanisms for shared information between subsets of modalities.
@@
To tackle the above issues, we propose a completed feature disentanglement multimodal learning (CFDL) approach for multimodal MRIs analysis. First, we present a novel completed feature disentanglement (CFD) strategy to address the information loss in previous FD-based methods. In addition to decoupling modality-shared features among all modalities and modality-specific features, We further decouple features shared between subsets of modalities, referred to as modality-partial-shared features. The modality partial-shared features are also expected to have higher similarities while being dissimilar from the other two kinds of features. As demonstrated in Section IV, these features play a critical role in prediction performance. Next, to improve the interpretability of feature fusion, we propose a new dynamic mixture-of-experts fusion (DMF) module, which can explicitly capture local-global interrelationships between the decoupled features to achieve more effective fusion. Finally, we evaluate our framework on three multimodal MRI datasets, demonstrating its effectiveness and superiority compared to state-of-the-art methods.
@@ 2.2 Multimodal Feature Fusion
In recent years, there has been an increasing focus on exploring the correlations among multiple modalities to obtain effective features. Some methods achieve multimodal fusion by assigning weights or probabilities to each modality. [15, 16, 17] explored the uncertainties of different modalities to obtain reliable multimodal fusion information. Choi et al. [21] proposed EmbraceNet which performed multimodal representation fusion based on a probabilistic approach. Zhou et al. [22] introduced a canonical correlation analysis (CCA)-based method named ADCCA to exploit the correlation between multiple modalities and integrate the complementary information from these modalities. Zhuang et al. [23] proposed a global-guided fusion method which consider both global and local correlations of multiple modalities.
@@ 3.1 Overview
The proposed CFDL framework, illustrated in Fig. 3, comprises three parts: a) feature extraction from multimodal MRIs, b) completed feature disentanglement for feature decoupling, and c) dynamic MoE fusion for dynamically integrating the decoupled features. The framework employs the same type of backbone to extract latent features $x_{j}$ from each modality. To capture a comprehensive representation of the multimodal data, the latent features are decoupled into modality-shared features, modality-specific features, and modality-partial-shared features using the proposed CFD strategy. The decoupled features are then integrated via the DMF module. Within the DMF module, each decoupled feature is paired with a specific expert, and a gating network named LinG_GN generates weights for multiple experts. The fused features are obtained by aggregating the weighted features from the experts. In the following, we take three-modal condition as examples to illustrate the CFD strategy and DMF module.
@@ 3.2 Completed Feature Disentanglement Strategy
Inspired by previous FD methods [7, 8, 9], we first decouple the extracted latent features into modality-shared features and modality-specific features. We employ a shared-encoder $\bm{E^{c}}$ to decouple modality-shared features and three private-encoders $\bm{E^{p}}_{j}$ to decouple modality-specific feature for each modality. Three modality-shared features can be formulated as follows:
$F_{j}=\bm{E^{c}}(x_{j}),$ (1)
@@
We further consider modality-partial-shared features between pair-wise modalities. As a result, three groups of modality-partial-shared features are decoupled, with each group consisting of two features. The two features in the same group ($G_{jk}^{j}$, $G_{jk}^{k}$) are decoupled with the same partial-shared-encoder named $\bm{E^{s}}_{jk}$, i.e.,
$G_{jk}^{j}=\bm{E^{s}}_{jk}(x_{j}),\quad G_{jk}^{k}=\bm{E^{s}}_{jk}(x_{k}).$ (4)
@@
To ensure the effective decoupling of modality-shared features and modality-partial-shared features, we employ the mean squared error (MSE) loss as a constraint. The MSE loss measures the discrepancy between two features, and we aim to increase similarity between two features by minimizing this loss. The losses for modality-shared features $\mathcal{L}_{sim}^{c}$ and modality-partial-shared features $\mathcal{L}_{sim}^{s}$ are expressed as:
$\mathcal{L}_{sim}^{c}=\sum_{j=1}\sum_{k=j+1}MSE(F_{j},F_{k}),$ (6)
$\mathcal{L}_{sim}^{s}=\sum_{j=1}\sum_{k=j+1}MSE(G_{jk}^{j},G_{jk}^{k}).$ (7)
We denote $\mathcal{L}_{sim}$ as the sum of $\mathcal{L}_{sim}^{c}$ and $\mathcal{L}_{sim}^{s}$.
@@ 3.3 Dynamic MoE Fusion Module
To ensure dynamic fusion of the final decoupled features, we introduce the DMF module based on the MoE architecture, which is shown in Fig. 3 (c). In DMF module, each final decoupled feature $\mathcal{S}_{j}\in\mathbb{R}^{dim}$ is associated with a specific expert $\bm{E^{x}}_{j}$, which is implemented as a fully-connected layer. We introduce LinG_GN to dynamically generate weights for these experts, taking into account the relationships of these final decoupled features. The LinG_GN operates with two inputs to capture a comprehensive understanding of these features. Firstly, we concatenate all the final decoupled features as one input for the LinG_GN. To integrate this concatenated feature, we map the concatenated feature into a unified representation utilizing a fully-connected layer $FC$. This unified representation captures the collective information from all the final decoupled features and is treated as the global feature $g\in\mathbb{R}^{dim}$,
@@
where $\times$ represents matrix multiplication operation, $\omega\in\mathbb{R}^{7}$. Each element in $\omega$ represents the weight for the corresponding expert network $\bm{E^{x}}_{j}$. The fused feature $F_{fused}$ is obtained by concatenating the linearly weighted experts,
$F_{fused}=Cat(\omega_{1}\bm{E^{x}}_{1}(\mathcal{S}_{1}),...,\omega_{7}\bm{E^{x}}_{7}(\mathcal{S}_{7})).$ (11)
@@
$\mathcal{L}=\mathcal{L}_{cls}+\alpha\mathcal{L}_{sim}+\beta\mathcal{L}_{diff},$ (13)
@@ 3.4 Network Architecture
We utilize the 3D ResNet18 [30] as the backbone for feature extraction, and the parameters of these backbones are not shared. The dimension of the each extracted latent feature $x_{j}$ is 512. The shared-encoder $\bm{E^{c}}$, each private-encoders $\bm{E^{p}}_{j}$, each partial-shared-encoders $\bm{E^{s}}_{jk}$, $FC$ and each expert $\bm{E^{x}}_{j}$ are all implemented as one fully-connected layer with $dim$ neurons. The $MLP$ consists of two fully-connected layers with $dim$ neurons and one output layer with $num\_cls$ neurons, where $num\_cls$ represents the number of classes. Each fully-connected layer in $MLP$ is followed by a ReLU layer and a Dropout layer. We empirically set $dim$ as 32.
@@ 4.1.1 Meniscal Tear Prediction
For the prediction of meniscal tear, we employ the MRNet dataset [31], which is a publicly available knee multi-MRI dataset [^1] More information about the MRNet dataset can be available in https://stanfordmlgroup.github.io/competitions/mrnet/.. There are 1130 cases in training set and 120 cases in validation set. The training set contains 397 meniscal tear cases and 733 contrast cases, and the validation set contains 52 meniscal tear cases and 68 contrast cases. Each case includes three MRIs: sagittal plane T2-weighted series (T2-sagittal), coronal plane T1-weighted series (T1-coronal) and axial plane PD-weighted series (PD-axial). We resize the MRIs to 24*128*128 as the model input.
@@ 4.1.2 Meningiomas Grading Prediction
We collect Meningiomas Grading Prediction Dataset, referred to as MEN, from the Brain Medical Center of Tianjin University, Tianjin Huanhu Hospital  [^2] The Ethical Committee of Tianjin Huanhu Hospital approves scientific research using these MRIs and waives the need for informed patient consent ((Jinhuan) Ethical Review No.(2022-046)).. This dataset consists of three grades of meningiomas: Grade 1 (G1), Grade 2 with invasion (G2inv) and Grade 2 without invasion (G2ninv). The total dataset comprises 798 cases, including 650 Grade_1, 62 Grade_2inv and 86 Grade_2ninv cases. Each case includes three brain MRIs: Contrast-Enhanced T1 series (T1C), Contrast-Enhanced T2 FLAIR series (FLAIR-C) and Apparent Diffusion Coefficient series (ADC). Following previous works [32, 33], we request radiologists to crop the regions of interest (ROIs). To maintain the shape of the tumor and edema regions, the ROIs are zero-padded into squares and resized to dimensions of 24*128*128, which serve as the inputs for the model.
@@ 4.1.3 MGMT Promoter Status Prediction
The MGMT Promoter Status Prediction Dataset, known as BraTS 2021 [34], is a publicly available multimodal brain MRI dataset. It encompasses cases with MGMT methylated (MGMT+) and unmethylated (MGMT-) status. The dataset comprises 580 available cases333 We drop 5 cases from original 585 cases during pre-processing. There are 3 cases with unexpected issues. The other 2 cases cannot be registered by CaPTk. Get more information in https://www.kaggle.com/c/rsna-miccai-brain-tumor-radiogenomic-classification., with each case containing four modalities: T1, post-contrast T1-weighted (T1Gd), T2-weighted (T2), and T2 Fluid Attenuated Inversion Recovery (T2-FLAIR). Specifically, there are 275 MGMT- cases and 305 MGMT+ cases. Pre-processing for each modality involves image registration and skull-stripping using the Cancer Imaging Phenomics Toolkit (CaPTk) [35]. We crop the ROIs using masks generated by the pretrained Swin UNETR [36]. Finally, we zero-pad the ROIs into squares and resize them to 16*128*128 as the inputs for the proposed method.
@@ 4.2.1 Training Details
We employ 3-fold cross-validation for private MEN and public BraTS 2021 datasets, and train three times using different seeds with already divided training and validation data for MRNet dataset. During model training, we implement several techniques to prevent overfitting, such as data augmentation, L2 regularization (weight decay) and dropout [37]. Data augmentation techniques include random clip, random crop, gaussian noise and random erasing [38]. The weight decay is set as $41e-$ 4, and the dropout value is set to $0.5$. The network is optimized with the Adam optimizer [39]. We linearly warm up the learning rate from zero to the preset value over $5$ epochs and apply a learning rate decay strategy, reducing the learning rate to $0.8$ after every $5$ epochs. The batch size is set $32$. For MRNet dataset, we initialize the learning rate value as $48e-$ 4, and set the number of epochs to $50$. For MEN dataset, the learning rate is specified as $45e-$ 4, and number of epochs is fixed as $100$. For BraTS 2021 dataset, we preset the learning rate value as $42e-$ 4 and the number of epochs as $50$. The balance factors, $\alpha$ and $\beta$, are set to {$0.5,0.005$}, {$1,1$}, {$0.001,0.001$} for MRNet, MEN and BraTS, respectively. Details of the ablation analysis for the balance factors are provided in the supplementary material. All experiments are conducted with PyTorch on an NVIDIA RTX 3090 GPU.